# AnyEnhance: A Unified Generative Model with Prompt-Guidance and Self-Critic for Voice Enhancement

### v1

@@ Abstract
We introduce AnyEnhance, a unified generative model for voice enhancement that processes both speech and singing voices. Based on a masked generative model, AnyEnhance is capable of handling both speech and singing voices, supporting a wide range of enhancement tasks including denoising, dereverberation, declipping, super-resolution, and target speaker extraction, all simultaneously and without fine-tuning. AnyEnhance introduces a prompt-guidance mechanism for in-context learning, which allows the model to natively accept a reference speaker’s timbre. In this way, it could boost enhancement performance when a reference audio is available and enable the target speaker extraction task without altering the underlying architecture. Moreover, we also introduce a self-critic mechanism into the generative process for masked generative models, yielding higher-quality outputs through iterative self-assessment and refinement. Extensive experiments on various enhancement tasks demonstrate AnyEnhance outperforms existing methods in terms of both objective metrics and subjective listening tests. Demo audios are publicly available at this URL.
Index Terms:
Generative model, speech enhancement, speech separation, target speaker extraction
@@ I Introduction
There are several models that aim to handle multiple enhancement tasks. Voicefixer [1] proposed the concept of General Speech Restoration (GSR), which attempts to remove multiple distortions simultaneously (noise, reverberation, clipping, and bandwidth limitation) from the input speech signal. MaskSR [2] is another GSR model using a masked generative model to handle noise, reverberation, clipping, and bandwidth limitation simultaneously. Although they can remove background noise from speech, they can’t separate speakers. SpeechX [3], SpeechFlow [4], NeMo [5] and UniAudio [6] are other multi-task enhancement models. SpeechX is a versatile speech generation model capable of tasks such as speech denoising and target speaker extraction. SpeechFlow and NeMo are both flow-matching-based [15] models that are pretrained on a large-scale dataset and can be fine-tuned for downstream tasks like speech denoising and target speaker extraction. UniAudio is an audio foundation model that supports 11 audio generation tasks covering speech denoising, dereverberation and target speaker extraction. Although those approaches can handle more speech enhancement tasks including target speaker extraction in one model, task-specific fine-tuning is still needed and they can’t handle multiple distortions in one pass (e.g., performing target speaker extraction tasks under noisy conditions). Also, they only work for speech enhancement, ignoring singing voice processing. Some works have attempted to address multi-domain enhancement problems. AudioSR [7] is a latent diffusion model which supports super-resolution across all audio domains (speech, sound effects, and music). But it only supports super-resolution tasks and does not include other enhancement tasks. Based on the above analysis, existing models have limitations in terms of both task and domain universality, and some require task-specific fine-tuning.
@@
To overcome the limitations, we propose a unified framework, AnyEnhance. Based on the masked generative model, AnyEnhance simultaneously addresses various enhancement tasks shown in Table I across both speech and singing voice domains, all without fine-tuning. To deal with real-world singing voice domain, we improved the data simulation process to create more realistic recording scenarios. Moreover, unlike implementing an auxiliary speaker encoder in many target speaker extraction models [10, 16], we propose a prompt-guidance mechanism in AnyEnhance, which provides the model with an in-context-learning capability. This mechanism enables the model to seamlessly accept a reference speaker’s voice for target speaker extraction and also improves performance on other tasks, all without architectural modifications. Furthermore, to address the issue of sampling instability inherent in masked generative models [17], we integrate a self-critic mechanism, enhancing accuracy and robustness when sampling from logits. These advancements make AnyEnhance a versatile solution for both speech and singing voice enhancement tasks, as well as a promising candidate for real-world applications across diverse domains. Experimental results demonstrate that AnyEnhance outperforms existing models across various enhancement tasks in terms of both objective metrics and subjective evaluations. Ablation studies further validate the effectiveness of the prompt-guidance mechanism across various tasks, the self-critic sampling strategy, the improved data simulation process, and the integration of various enhancement tasks into a unified model. To summarize, our contributions are as follows:
@@ III-A Enhancement Problem Formulation
  • Bandwidth Limitation: $d_{\text{bw}}(x)=Resample(x,freq)$, where $freq$ is the target lower sampling frequency.
@@
When simulating data, previous works always have the order of $d_{\text{noise}}(d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(\cdot))))$ which puts $d_{\text{noise}}$ at the final stage. However, in practice, noise would be intertwined with other distortions (like room reverb). Therefore, to achieve a more realistic data simulation, we propose an improved data simulation strategy where noise is added before reverberation and $d_{\text{bw}}$ and $d_{\text{clip}}$ are placed at the end of the distortion chain. We also add a separate eq-reverberation chain before adding noise to simulate real-world live performance vocal effects. The final distortion chain is:
@@
where $d_{\text{vocal\_effect}}=d_{\text{reverb}}(d_{\text{eq}}(x))$ represents the vocal effect chain. Here, $d_{\text{eq}}$ denotes the equalization effect, which applies 1-3 bell gains in the range of [-5, 5] dB across 10-12,000 Hz within a 1-second window. We define two general tasks for our model: Enhancement and Extraction. The enhancement task aims to recover the clean audio from the distorted audio, while the extraction task aims to extract the target speaker’s voice from the mixture. The difference between these two tasks is that the extraction task would insert $d_{\text{voice}}$ into the head of the distortion chain.
@@ III-B Masked Generative Modeling
[…]
@@ IV-A Model Architecture
During the semantic enhancement stage, the input of the encoder is the power-law compressed magnitude STFT spectrogram [30] of the distorted audio $D(x)$ and task embedding $h_{\text{task}}$. The encoder consists of $L_{enc}$ self-attention transformer layers to obtain the latent representation $h_{\text{enc}}\in\mathbb{R}^{T\times dim}$. Inspired by representation alignment (REPA) in the field of image generation [31] and semantic distillation in MaskSR2 [26], we introduce a representation alignment module to align the encoder output. The semantic feature of the clean audio is extracted using self-supervised learning (SSL) models, such as w2v-BERT [32]. Let $f$ be a SSL encoder, $z=f(x)\in\mathbb{R}^{T\times dim}$ be an encoder output (linear interpolation is used to match time dimension $T$ and feature dimension $dim$), a multilayer perceptron $\text{MLP}_{\phi}$ is used to project the encoder output with the clean audio’s semantic feature:
@@
where the similarity function is mean-square error (MSE). The final output of the semantic enhancement stage is the sum of encoder’s output and noisy semantic feature $h_{\text{semantic}}=h_{\text{enc}}+f(D(x))$.
@@ IV-B Prompt Guidance
We introduce a prompt-guidance scheme (see Figure 2). The model can be trained using both configurations: without prompt and with prompt. During training, we have a probability $p_{\text{prompt}}$ to keep the first 3 seconds of clean audio unmasked as the prompt. The same place of the distorted audio is set to silence. The prompt section is not considered for loss calculation. During inference, the model can infer both with a prompt audio and without prompt audio:
@@ V-A Datasets
Training Datasets The training datasets used in our experiments include speech, singing voice, noise, and room impulse responses (RIRs), with detailed statistics shown in Table II. For speech data, we used 24kHz speech from Emilia [14] Subset (filtered with DNSMOS OVRL $>$ 3.4, totaling $\sim$21k hours) and fullband speech from datasets such as VCTK [35], HiFi-TTS [36] ($\sim$1k hours), and AI-Shell3 [37]. Singing voice data include OpenSinger [39], PopCS [40], and others ($\sim$600 hours). Noise data are sourced from the musan noise subset [46], Urgent Challenge noise data [21], FSD50K [47], DESED [48], and the TUT Urban Acoustic Scenes 2018 dataset [49], along with a proprietary dataset of live-performance noise ($\sim$1.2k hours). RIRs comprise 62,668 pieces from SLR26, SLR28, and self-collected RIRs from various environments. We dynamically distort the voice with noise and RIR data to get the training data pair for the Enhancement and Extraction tasks described in Section III-A. All the audio pieces are resampled to 44.1kHz. Table III shows the probability and hyperparameters of the distortion used for training data.
@@ V-C Implementation Details
We use the pretrained 44.1kHz DAC [25] codec as our audio tokenizer. For the main experiments, we use a configuration of $dim=768,L_{enc}=24,L_{dec}=24$ as configured in Section IV-A (768 hidden dim, 24 transformer layers for Encoder, 24 transformer layers for Decoder in Figure 1), totaling of 363.54M parameters, designed to comprehensively evaluate the model’s enhancement capabilities for both speech and singing voice, utilizing the full dataset and the comprehensive data degradation pipeline. We use rotary position embedding RoPE [58] for the transformer layers as observed in [59]. We employ the 17th layer of w2v-bert2111https://huggingface.co/facebook/w2v-bert-2.0 for semantic feature alignment. Each audio sample is 8 seconds long, with a 50% probability of retaining the first 3 seconds as the prompt. Training is conducted on 4 A800 GPUs, with a batch size of 8 per GPU using Adam optimizer for 300k steps, with a learning rate of 1e-4, a 4k-step warmup and linear decay to 0.
For the task and distortion chain ablation studies, we use a smaller model with $dim=512,L_{enc}=8,L_{dec}=12$, totaling 77.93M parameters. The model is trained on the 44.1kHz speech dataset and a subset of noise data totaling 400 hours. Each audio sample is 6 seconds long. Training is conducted on 2 A800 GPUs, with a batch size of 16 per GPU. The rest of the settings remain the same as in the main experiment.
@@ V-D Evaluation Metrics
Some common signal-based metrics, such as PESQ [63], STOI [64], and ESTOI [65], are often used to evaluate enhancement models. However, these metrics are unsuitable for generative-model-based enhancement models due to the lack of waveform-level alignment between the enhanced and reference signals [2, 22, 16]. As a result, we do not include these metrics in our evaluations. Instead, we focus on metrics that better capture the perceptual and semantic qualities of the enhanced audio.
@@ V-E1 Evaluation on Various Enhancement Tasks
V-E2 Effect of Prompt Guidance
TABLE VII: Effect of Prompt Guidance, AnyEnhance can achieve better performance with a speaker prompt.

### v3

@@ Abstract
We introduce AnyEnhance, a unified generative model for voice enhancement that processes both speech and singing voices. Based on a masked generative model, AnyEnhance is capable of handling both speech and singing voices, supporting a wide range of enhancement tasks including denoising, dereverberation, declipping, super-resolution, and target speaker extraction, all simultaneously and without fine-tuning. AnyEnhance introduces a prompt-guidance mechanism for in-context learning, which allows the model to natively accept a reference speaker’s timbre. In this way, it could boost enhancement performance when a reference audio is available and enable the target speaker extraction task without altering the underlying architecture. Moreover, we also introduce a self-critic mechanism into the generative process for masked generative models, yielding higher-quality outputs through iterative self-assessment and refinement. Extensive experiments on various enhancement tasks demonstrate AnyEnhance outperforms existing methods in terms of both objective metrics and subjective listening tests. Demo audios are publicly available at  [^1] https://amphionspace.github.io/anyenhance. An open-source implementation is provided at  [^2] https://github.com/viewfinder-annn/anyenhance-v1-ccf-aatc.
@@ I Introduction
There are several models that aim to handle multiple enhancement tasks. Voicefixer [1] proposed the concept of General Speech Restoration (GSR), which attempts to remove multiple distortions simultaneously (noise, reverberation, clipping, and bandwidth limitation) from the input speech signal. MaskSR [2] is another GSR model using a masked generative model to handle noise, reverberation, clipping, and bandwidth limitation simultaneously. Although they can remove background noise from speech, they can’t separate speakers. SpeechX [3], SpeechFlow [4], NeMo [5] and UniAudio [6] are other multi-task enhancement models. SpeechX is a versatile speech generation model capable of tasks such as speech denoising and target speaker extraction. SpeechFlow and NeMo are both flow-matching-based [15] models that are pretrained on a large-scale dataset and can be fine-tuned for downstream tasks like speech denoising and target speaker extraction. UniAudio is an audio foundation model that supports 11 audio generation tasks covering speech denoising, dereverberation and target speaker extraction. Although those approaches can handle more speech enhancement tasks including target speaker extraction in one model, task-specific fine-tuning is still needed and they can’t handle multiple distortions in one pass (e.g., performing target speaker extraction tasks under noisy conditions). Also, they only work for speech enhancement, ignoring singing voice processing. Some works have attempted to address multi-domain enhancement problems. AudioSR [7] is a latent diffusion model which supports super-resolution across all audio domains (speech, sound effects, and music), but it only supports super-resolution tasks and does not include other enhancement tasks. Based on the above analysis, existing models have limitations in terms of both task and domain universality, and some require task-specific fine-tuning.
@@
To overcome the limitations, we propose a unified framework, AnyEnhance. Based on the masked generative model, AnyEnhance simultaneously addresses various enhancement tasks shown in Table I across both speech and singing voice domains, all without fine-tuning. To deal with real-world singing voice domain, we improved the data simulation process to create more realistic recording scenarios. Moreover, unlike implementing an auxiliary speaker encoder in many target speaker extraction models [10, 16], we propose a prompt-guidance mechanism in AnyEnhance, which provides the model with an in-context-learning capability. This mechanism enables the model to seamlessly accept a reference speaker’s voice for target speaker extraction and also improves performance on other tasks, all without architectural modifications. Furthermore, to address the issue of sampling instability inherent in masked generative models [17], we integrate a self-critic mechanism, enhancing accuracy and robustness when sampling from logits. These advancements make AnyEnhance a versatile solution for both speech and singing voice enhancement tasks, as well as a promising candidate for real-world applications across diverse domains. Experimental results demonstrate that AnyEnhance outperforms existing models across various enhancement tasks in terms of both objective metrics and subjective evaluations. Ablation studies further validate the effectiveness of the prompt-guidance mechanism across various tasks, the self-critic sampling strategy, the improved data simulation process, and the integration of various enhancement tasks into a unified model. Our unified multitask training consistently outperforms single-task enhancement models across all tasks, with average improvements of +0.067 OVRL and +0.13 NISQA, demonstrating strong generalization across diverse speech and singing tasks. To summarize, our contributions are as follows:
@@ III-A Enhancement Problem Formulation
  • Bandwidth Limitation: $d_{\text{bw}}(x)=Resample(x,freq)$, where freqfreq is the target lower sampling frequency.
@@
Figure 2: Simulation of real-world live vocal recordings, including vocal effects, venue acoustics, and ambient noise.
When simulating data, previous works typically follow the distortion order of $d_{\text{noise}}(d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(\cdot))))$, placing additive noise at the final stage. However, in real-world scenarios, noise is often intertwined with reverberation and other artifacts. An overview of the real-world recording path simulated in our framework is illustrated in Figure 2. To better approximate practical conditions, we propose a refined data simulation strategy that applies noise before reverberation, and places bandwidth and clipping effects at the end. Additionally, we insert a separate vocal effects chain (e.g., EQ, reverb) prior to noise to simulate live performance processing. The final degradation chain is formulated as:
@@
where $d_{\text{vocal\_effect}}=d_{\text{reverb}}(d_{\text{eq}}(x))$ represents the vocal effect chain. Here, $d_{\text{eq}}$ denotes the equalization effect, which applies 1-3 bell gains in the range of [-5, 5] dB across 10-12,000 Hz within a 1-second window. We define two general tasks for our model: Enhancement and Extraction. The enhancement task aims to recover the clean audio from the distorted audio, while the extraction task aims to extract the target speaker’s voice from the mixture. The difference between these two tasks is that the extraction task would insert $d_{\text{voice}}$ into the head of the distortion chain, namely $D_{extraction}(x)=d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(d_{\text{voice}}(d_{\text{vocal\_effect}}(x))))).$
@@ III-B Masked Generative Modeling
During inference, the reverse diffusion process iteratively predicts the tokens starting from fully masked token sequence $\mathbf{X}_{T}$. Suppose the total number of inference steps is $S$, at each timestep $t=\{S,S-1,\dots,1\}$, we first sample $\hat{\mathbf{X}}_{0}$ from the model’s logits $p_{\theta}(\mathbf{X}_{0}\mid\mathbf{X}_{\frac{t}{S}T},\mathbf{C})$. Then, we choose $\lfloor N\cdot\gamma(\frac{t-1}{S}T)\rfloor$ tokens to remask based on the confidence score. In common practice [28, 2, 29], the model’s logits $p_{\theta}(\hat{x}_{i}\mid\mathbf{X}_{\frac{t}{S}T},\mathbf{C})$ is directly taken as confidence score if $\mathbf{X}_{\frac{t}{S}T,i}=\text{[MASK]}$, otherwise 1, and tokens with the lowest confidence scores are remasked.
@@ IV-A Model Architecture
During the semantic enhancement stage, the input of the encoder is the power-law compressed magnitude STFT spectrogram [30] of the distorted audio $D(x)$ and task embedding $h_{\text{task}}$. The encoder consists of $L_{enc}$ self-attention transformer layers to obtain the latent representation $h_{\text{enc}}\in\mathbb{R}^{T\times dim}$. Inspired by representation alignment (REPA) in the field of image generation [31] and semantic distillation in MaskSR2 [26], we introduce a representation alignment module to align the encoder output. The semantic feature of the clean audio is extracted using self-supervised learning (SSL) models, such as w2v-BERT [32]. Let $f$ be a SSL encoder, $z=f(x)\in\mathbb{R}^{T\times dim}$ be an encoder output (linear interpolation is used to match time dimension $T$ and feature dimension dimdim), a multilayer perceptron $\text{MLP}_{\phi}$ is used to project the encoder output with the clean audio’s semantic feature:
@@
where the similarity function is mean-square error (MSE). Furthermore, motivated by [33, 34], we also integrate self-supervised feature to the output for semantic enhancement. Prior work has shown that SSL models possess inherent noise robustness and can significantly improve speech enhancement performance when used as input features. Based on this insight, the final output of our semantic enhancement stage is formulated as the sum of the encoder output and the noisy semantic feature: $h_{\text{semantic}}=h_{\text{enc}}+f(D(x))$.
@@
Though AnyEnhance adopts MaskSR’s architecture, it introduces several key innovations, including degradation path optimization, prompt-guidance and self-critic sampling. These concepts are generally applicable and beneficial beyond the MaskSR framework. While MaskSR focuses solely on speech restoration, AnyEnhance extends its applicability to multiple tasks such as target speech extraction and singing voice enhancement. Our prompt-guidance mechanism is not limited to enabling the TSE task. Inspired by the timbre cloning capability in zero-shot TTS, we aim to leverage prompt guidance to enhance timbre similarity and overall quality across a broader range of tasks.
@@ IV-B Prompt Guidance
We introduce a prompt-guidance scheme (see Figure 3). The design of the prompt-guidance mechanism is inspired by the success of zero-shot TTS systems [35, 29], where models can mimic the timbre and audio characteristics of a reference speaker through prompt conditioning. Our goal is to extend this capability to speech enhancement tasks, enabling the model to leverage clean prompt audio to learn target speaker characteristics and audio quality in an end-to-end manner without alternating the model architecture. The model can be trained using both configurations: without prompt and with prompt. During training, we have a probability $p_{\text{prompt}}$ to keep the first 3 seconds of clean audio unmasked as the prompt. The same place of the distorted audio is set to silence. The prompt section is not considered for loss calculation. During inference, the model can infer both with a prompt audio and without prompt audio:
@@ V-A Datasets
Training Datasets The training datasets used in our experiments include speech, singing voice, noise, and room impulse responses (RIRs), with detailed statistics shown in Table II. For speech data, we used 24kHz speech from Emilia [14] Subset (filtered with DNSMOS OVRL >> 3.4, totaling $\sim$21k hours) and fullband speech from datasets such as VCTK [38], HiFi-TTS [39] ($\sim$1k hours), and AI-Shell3 [40]. Singing voice data include OpenSinger [42], PopCS [43], and others ($\sim$600 hours). Noise data are sourced from the musan noise subset [49], Urgent Challenge noise data [21], FSD50K [50], DESED [51], and the TUT Urban Acoustic Scenes 2018 dataset [52], along with a proprietary dataset of live-performance noise ($\sim$1.2k hours). RIRs comprise 62,668 pieces from SLR26, SLR28, and self-collected RIRs from various environments. We dynamically distort the voice with noise and RIR data to get the training data pair for the Enhancement and Extraction tasks described in Section III-A. All the audio pieces are resampled to 44.1kHz. Table III shows the probability and hyperparameters of the distortion used for training data.
@@ V-C Implementation Details
We use the pretrained 44.1kHz DAC [25] codec as our audio tokenizer. For the main experiments, we use a configuration of $dim=768,L_{enc}=24,L_{dec}=24$ as configured in Section IV-A (768 hidden dim, 24 transformer layers for Encoder, 24 transformer layers for Decoder in Figure 1), totaling of 363.54M parameters, designed to comprehensively evaluate the model’s enhancement capabilities for both speech and singing voice, utilizing the full dataset and the comprehensive data degradation pipeline. We use rotary position embedding RoPE [61] for the transformer layers as observed in [62]. We employ the 17th layer of w2v-bert2333https://huggingface.co/facebook/w2v-bert-2.0 for semantic feature alignment. Each audio sample is 8 seconds long, with a 50% probability of retaining the first 3 seconds as the prompt. Training is conducted on 4 GPUs, with a batch size of 8 per GPU using Adam optimizer for 300k steps, with a learning rate of 1e-4, a 4k-step warmup and linear decay to 0.
For the task and distortion chain ablation studies, we use a smaller model with $dim=512,L_{enc}=8,L_{dec}=12$, totaling 77.93M parameters. The model is trained on the 44.1kHz speech dataset and a subset of noise data totaling 400 hours. Each audio sample is 6 seconds long. Training is conducted on 2 GPUs, with a batch size of 16 per GPU. The rest of the settings remain the same as in the main experiment.
@@ V-D Evaluation Metrics
Some common signal-based metrics, such as PESQ [66], STOI [67], and ESTOI [68], are often used to evaluate enhancement models. However, these metrics are unsuitable for generative-model-based enhancement models due to the lack of waveform-level alignment between the enhanced and reference signals [2, 22, 16]. As a result, we do not include these metrics in our main evaluations and instead focus on perceptual and semantic-level metrics that better reflect the quality of generative outputs. Nevertheless, for a comprehensive evaluation and analysis, we still report these alignment-dependent metrics, as well as the Word Error Rate (WER), in Table VII for the GSR Group.
@@ V-E1 Evaluation on Various Enhancement Tasks
Additionally, we report the intelligibility metric (Word Error Rate, WER) and signal-based metrics (PESQ, ESTOI) on the Voicefixer GSR, Librivox GSR, and CCMusic GSR datasets in Table VII. We use Whisper-large-v3666https://huggingface.co/openai/whisper-large-v3, a state-of-the-art ASR model, for WER evaluation. However, it sometimes fails to transcribe even the ground-truth audios in the CCMusic GSR test set, resulting in unreliable WER scores. So we do not report WER for this dataset. The results show that AnyEnhance consistently achieves competitive or superior performance across all three datasets. On Voicefixer GSR AnyEnhance ranks second in WER, PESQ, and ESTOI, close behind TFGridNet. On Librivox GSR, AnyEnhance yields the lowest WER (0.199), the second-highest PESQ (2.073), and the best ESTOI (0.784). On the singing voice domain (CCMusic GSR), AnyEnhance significantly surpasses all other systems in both PESQ (2.337) and ESTOI (0.585), showing its strong generalization ability in singing-voice-oriented enhancement scenarios.
TABLE VII: Intelligibility and signal-based metric comparison (WER, PESQ, ESTOI) on Voicefixer GSR, Librivox GSR, and CCMusic GSR test sets. WER is not reported for CCMusic GSR due to ASR transcription limitations.
[…]
Voicefixer GSR TFGridNet 0.092 2.47 0.672
NSNet2 0.149 2.09 0.608
Voicefixer (checkpoint) 0.141 2.034 0.557
Voicefixer (retrained) 0.134 2.001 0.563
MaskSR 0.147 2.319 0.658
AnyEnhance 0.113 2.449 0.667
Librivox GSR TFGridNet 0.201 2.076 0.708
NSNet2 0.327 1.744 0.619
Voicefixer (checkpoint) 0.363 1.561 0.622
Voicefixer (retrained) 0.24 1.588 0.663
MaskSR 0.262 1.859 0.743
AnyEnhance 0.199 2.073 0.784
CCMusic GSR TFGridNet - 1.417 0.4
NSNet2 - 1.474 0.396
Voicefixer (checkpoint) - 1.464 0.399
Voicefixer (retrained) - 1.66 0.476
MaskSR - 2.1 0.541
AnyEnhance - 2.337 0.585
TABLE VIII: Evaluation on real-world recordings from the DNS Challenge dataset.
[…]
DNS Real Recordings TFGridNet 3.321 3.844 2.973
NSNet2 3.037 3.963 2.76
Voicefixer (checkpoint) 3.289 3.955 2.988
Voicefixer (retrained) 3.218 3.892 2.9
MaskSR 3.33 4.044 3.056
AnyEnhance 3.488 3.977 3.161
We also evaluate the performance of AnyEnhance on the DNS Challenge Real Recordings dataset, which contains real-world recordings. The results are shown in Table VIII. AnyEnhance achieves the highest SIG and OVRL scores among all models, indicating strong real-world enhancement capability, while maintaining competitive BAK performance.
TABLE IX: Training time, model size, and average real-time factor (RTF) comparison on the DNS No Reverb dataset. The RTF is measured on a single GPU. Lower RTF indicates faster inference.
[…]
Voicefixer (retrained) 111.7 M 87.222 h 0.105
MaskSR 361.8 M 64.118 h 0.259
AnyEnhance 363.6 M 125.833 h 0.254
We report the model size, training time, and average real-time factor (RTF) of Voicefixer, MaskSR, and AnyEnhance in Table IX. All models are trained on the same dataset using 4 GPUs, and the RTF is evaluated on the DNS No Reverb dataset using a single GPU. AnyEnhance requires the longest training time (125.8 hours) due to its self-critic and REPA design, but achieves competitive inference efficiency (RTF = 0.254) comparable to MaskSR (0.259). While Voicefixer achieves the fastest inference (RTF = 0.105), all models operate well below the real-time threshold (RTF << 1), demonstrating their practical efficiency for deployment.
V-E2 Effect of Prompt Guidance and Self-Critic Sampling
TABLE X: Effect of Prompt Guidance, AnyEnhance can achieve better performance with a speaker prompt.